OpenStudy (anonymous):

Solve the rational equation by using the LCD. (x)/(x^2-16)+(4)/(x-4)=(1)/(x+4)

OpenStudy (anonymous):

hint factor out the (x^2-16)

OpenStudy (anonymous):

x=-5?

OpenStudy (anonymous):

where did u get the x=-5

OpenStudy (anonymous):

(5x+16)/((x-4)(x+4))=(1)/(x+4)

OpenStudy (anonymous):

did u cross multiply?

OpenStudy (anonymous):

yes, then I got (16)/((x-4)(x+4))=(1)/(x+4) = x-20/x-4 = -5

OpenStudy (anonymous):

 this what i got